at the close of one business day, a company's stock was trading at $24 a share. at the close of the next business day, the stock was trading at 27$ a share. find the percent of increase.
Answer by Alan3354(69443)

at the close of one business day, a company's stock was trading at $24 a share. at the close of the next business day, the stock was trading at 27$ a share. find the percent of increase.
----------------
It increased $3.
3/24 = 1/8
100% times 1/8 = 12.5%.
That's the increase, 12.5%